E10.13 (LO 1, 3) (Entries for Acquisition of Assets) Presented below is information related to Zonker Company.

1.   On July 6, Zonker Company acquired the plant assets of Doonesbury Company, which had discontinued operations. The appraised value of the property is:

Land
Buildings
Equipment

$400,000
1,200,000
800,000

    Total

$2,400,000

   Zonker Company gave 12,500 shares of its $100 par value common stock in exchange. The stock had a market price of $168 per share on the date of the purchase of the property.

How do I calculate how much each asset is to receive? Example 2100000 (12500*168). But I'm not sure how to assign.
Answer #1

SOLUTION:

Cost of the property, plant and equipment: $12,500 * $168 i= $2,100,000

Allocated cost is based on appraisal values, thus:

Land: 400,000 / 2,400,000 * 2,100,000 = $350,000

Building: 1,200,000 / 2,400,000 * 2,100,000 = $1,050,000

Equipment: 800,000 / 2,400,000 * 2,100,000 = $700,000

 

Journal entry:

Particulars

Debit

Credit

Land

350,000

Building

1,050,000

Equipment

700,000

Common stock (12,500 * $100)

1,250,000

Paid-in Capital in Excess of Par - Common Stock

850,000

> Thanks for the correct answer and explanation!

I'minaccountinghell Sat, Apr 8, 2023 8:32 PM